# Weak instrumental variables due to nonlinearities in panel data: A Super Learner Control Function estimator

### v6

@@ Abstract
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders with some of them time-invariant. In this setup, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to weak instruments (instruments are weakly correlated with the endogeneous covariate). As a solution, we propose a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. First, we estimate the control function using a super learner. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. The estimator is consistent and asymptotically normal achieving a parametric rate of convergence. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the Super Learner Control Function Estimators significantly outperform Within 2SLS estimators. Finally, we show that the SLCF estimator is not numerically equivalent to a naive plugin IV estimator.
@@ 1 Introduction
In Section 3, we present motivating examples. In Section 4, we present the baseline model. In Section 5 we describe the identification strategy. In Section 6 we present the SLCF estimator. In Section 8 we present the large sample properties of the SLCF estimator, in Section 10, we present the simulation experiment and its results, in Section 11 we present the results of the empirical application, and in Section 12 we present the conclusions of the paper.
@@ 8.1.1 Including sample splitting uncertainty
9 Naive plugin IV estimation is inconsistent and it is not numerically equivalent to SLCFE
The Super Learner Control function estimator is not numerical equivalent to an IV estimator that plugs in the estimated transformed nuisance parameter $\tau g_{o}(\tilde{x}_{it},z_{it})$. Without loss of generality, we assume that exogenous covariates are not present (model 1 to 3). As described in Section 5, one could be tempted to estimate the parameter of insterest by using the naive IV estimator:
@@ 9 Naive plugin IV estimation is inconsistent and it is not numerically equivalent to SLCFE
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, Naive plugin IV estimation is not numerically equivalent to SLCFE when $g_{o}(.)$ is nonlinear
Proof 2.
@@
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold in particular that $\left\|\tau g-\tau g_{o}\right\|_{P,2}\leq\delta_{N}$ with $\delta_{N}\sqrt{N_{T}}\to\infty$, $\{\delta_{N}\}$ a sequence of positive numbers that converge to 0 as $N\rightarrow\infty$ , the Naive plugin IV estimator is inconsistent.
Proof 3.
The naive plugin IV estimator with sample splitting is given by:
$\hat{\beta}_{IV,1_{o}}=\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}\Big).$ (41)
$\hat{\beta}_{1,o}=\beta_{1,o}+\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}(\xi_{i}^{S_{b}^{c}}\beta_{1,o}+M_{\tau_{i}}u_{i}+M_{\tau_{i}}\varepsilon_{i})\Big),$ (42)
with $\xi_{i}^{S_{b}^{c}}=\tau g_{o}(z_{i})-\widehat{\tau g_{o}}(z_{i})^{S_{b}^{c}}$.
$\hat{\beta}_{1,o}=\beta_{1,o}+\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i\in S^{b}}(\xi^{S_{b}^{c}}_{i}+\tau g_{o}(z_{i}))^{\prime}(\xi^{S_{b}^{c}}_{i}\beta_{1,o}+M_{\tau_{i}}u_{i}+M_{\tau_{i}}\varepsilon_{i})\Big),$ (43)
The second term does not converge to zero at rate faster than $\sqrt{N_{T}}$ because $\tau g_{o}(z_{i})$ is not centered at 0. Thus, the naive plugin estimator presents a first order bias term that does not vanish using sample-splitting. Thus, the naive plugin IV estimator is inconsistent.

### v7

@@ Abstract
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders with some of them time-invariant. In this setup, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to weak instruments (instruments are weakly correlated with the endogeneous covariate). As a solution, we propose a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. First, we estimate the control function using a super learner. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. The estimator is consistent and asymptotically normal achieving a parametric rate of convergence. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the Super Learner Control Function Estimators significantly outperform Within 2SLS estimators. Finally, we show that the SLCF estimator differs from both the plug-in IV estimator and a naive plug-in 2SLS estimator, with the former being incosistent without cross-fitting, and the latter being inconsitent even with cross-fitting.
@@ 1 Introduction
In Section 3, we present motivating examples. In Section 4, we present the baseline model. In Section 5, we describe the identification strategy. In Section 6, we present the SLCF estimator. In Section 8, we present the large sample properties of the SLCF estimator. In Section 9, we present a comparison of the proposed estimator with IV and 2SLS plugin estimators. In Section 10, we present the simulation experiment and its results, in Section 11 we present the results of the empirical application, and in Section 12 we present the conclusions of the paper.
@@ 8.1.1 Including sample splitting uncertainty
9 Comparison of the SLCFE with plug-in IV and naive plug-in 2SLS estimators
9.1 Naive plug-in 2SLS estimator is inconsistent and it is not numerically equivalent to SLCFE
The Super Learner Control function estimator is not numerical equivalent to a 2SLS estimator that plugs in the estimated transformed nuisance parameter $\tau g_{o}(\tilde{x}_{it},z_{it})$. Without loss of generality, we assume that exogenous covariates are not present (model 1 to 3). As described in Section 5, one could be tempted to estimate the parameter of insterest by using the naive 2SLS estimator:
$\hat{\beta}_{2SLS,1_{o}}=\sum_{b}^{B}\frac{1}{B}(\sum_{i\in S_{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i}))^{-1}(\sum_{i\in S_{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}).$ (38)
Theorem 4.
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, Naive plug-in 2SLS estimation is not numerically equivalent to SLCFE when $g_{o}(.)$ is nonlinear
Proof 2.
Assuming that $B=1$, we can write the estimators as:
$\hat{\beta}_{2SLS,1_{o}}=\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})\Big)^{-1}\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}\Big),$ (39)
Theorem 5.
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold in particular that $\left\|\tau g-\tau g_{o}\right\|_{P,2}\leq\delta_{N}$ with $\delta_{N}\sqrt{N_{T}}\to\infty$, $\{\delta_{N}\}$ a sequence of positive numbers that converge to 0 as $N\rightarrow\infty$ , the Naive plug-in 2SLS estimator is inconsistent.
Proof 3.
The naive plug-in 2SLS estimator with sample splitting is given by:
$\hat{\beta}_{2SLS,1_{o}}=\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})\Big)^{-1}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}\Big).$ (41)
$\hat{\beta}_{2SLS,1_{o}}=\beta_{1_{o}}+\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})\Big)^{-1}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}(\xi_{i}^{S_{b}^{c}}\beta_{1,o}+M_{\tau_{i}}u_{i}+M_{\tau_{i}}\varepsilon_{i})\Big),$ (42)
with $\xi_{i}^{S_{b}^{c}}=\tau g_{o}(z_{i})-\widehat{\tau g_{o}}(z_{i})^{S_{b}^{c}}$.
$\hat{\beta}_{2SLS,1_{o}}=\beta_{1_{o}}+\frac{1}{B}\sum_{b}^{B}\Big(\sum_{i\in S^{b}}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})^{\prime}\widehat{\tau g_{o}}^{S_{b}^{c}}(z_{i})\Big)^{-1}\Big(\sum_{i\in S^{b}}(\xi^{S_{b}^{c}}_{i}+\tau g_{o}(z_{i}))^{\prime}(\xi^{S_{b}^{c}}_{i}\beta_{1,o}+M_{\tau_{i}}u_{i}+M_{\tau_{i}}\varepsilon_{i})\Big),$ (43)
The second term does not converge to zero at rate faster than $\sqrt{N_{T}}$ because $\tau g_{o}(z_{i})$ is not centered at 0. Thus, the naive plug-in estimator presents a first order bias term that does not vanish using sample-splitting. Thus, the naive plug-in 2SLS estimator is inconsistent.
9.2 Plug-in IV estimation is not numerically equivalent to SLCFE
The IV estimator is given by:
@@ 9.2 Plug-in IV estimation is not numerically equivalent to SLCFE
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold, plug-in IV estimation is not numerically equivalent to SLCFE when $g_{o}(.)$ is nonlinear
Proof 4.
@@
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 13.3 in the Annex hold in particular that $\left\|\tau g-\tau g_{o}\right\|_{P,2}\leq\delta_{N}$ with $\delta_{N}\sqrt{N_{T}}\to\infty$, $\{\delta_{N}\}$ a sequence of positive numbers that converge to 0 as $N\rightarrow\infty$ , the plug-in IV estimator without the use of sample splitting is inconsistent.
Proof 5.
The plug-in IV estimator without sample splitting is given by:
$\hat{\beta}_{IV,1_{o}}=\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}y_{i}\Big).$ (47)
Replacing $M_{\tau_{i}}y_{i}$ by $M_{\tau_{i}}x_{1i}\beta_{1_{o}}+M_{\tau_{i}}\varepsilon_{i}$, it can be written as:
$\hat{\beta}_{IV,1_{o}}=\beta_{1_{o}}+\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}(M_{\tau_{i}}\varepsilon_{i})\Big),$ (48)
Now, adding and substracting $\tau g_{o}(z_{i})$ to $\widehat{\tau g_{o}}(z_{i})$ we obtain that:
$\hat{\beta}_{IV,1_{o}}=\beta_{1_{o}}+\Big(\sum_{i}\widehat{\tau g_{o}}(z_{i})^{\prime}M_{\tau_{i}}x_{1i}\Big)^{-1}\Big(\sum_{i}(\xi_{i}+\tau g_{o}(z_{i}))^{\prime}(M_{\tau_{i}}\varepsilon_{i})\Big),$ (49)
The second term does not converge to zero at rate faster than $\sqrt{N_{T}}$ because $\tau g_{o}(z_{i})$ is not learned using sample splitting. Thus, the plug-in IV estimator without using sample-splitting presents a first order bias term that does not vanish. Thus, the plug-in IV estimator is inconsistent when estimation of the transformed nuisance parameter is done without cross-fitting.